OpenStudy (anonymous):

Use a matrix to find a solution to the system of equations: -8x-8y=-16 6x-9y=-108 Answers: (-6,8) (6,8) (8,-6) (6,-8)

OpenStudy (anonymous):

Please explain! I'm really stuck, my teacher didn't explain matrices well and I have no idea how to solve this....

OpenStudy (anonymous):

@Hero please help!

hero (hero):

I saw this exact same question before and I answered it as well.

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

well can you please help me? <3

hero (hero):

Okay, I know how to solve it using Gaussian Elimination method

hero (hero):

Which do you use.. Gaussian Elimination Ax = B or Cramer's Rule?

OpenStudy (anonymous):

Um I have no idea what either of those are....my teacher is kind of bad

OpenStudy (anonymous):

I know what eliminations is, but I've never heard of Gaussian Elimination

hero (hero):

Okay, what about Reduced Row Echelon Form?

OpenStudy (anonymous):

yeah, that's where it ends up being all 1s and 0s with the constant on the right side, correct?

hero (hero):

Yes, good. That's the one that I am comfortable with because it is easy.

OpenStudy (anonymous):

okay, that works for me

hero (hero):

So basically, you have this \[\left[ \begin{matrix} -8 & -8 & -16 \\ 6 & 9 & 108 \end{matrix} \right] \]

OpenStudy (anonymous):

yes

hero (hero):

The first thing you can do is divide the first row by -8 and the second row by 3. When you do that you will get: \[\left[ \begin{matrix} 1 & 1 & 2 \\ 2 & 3 & 36 \end{matrix} \right]\]

hero (hero):

Now from there, you can multiply the 1st row by 2, then subtract Row 2 - 2*Row 1: \[\left[ \begin{matrix} 1 & 1 & 2 \\ 0 & 1 & 32 \end{matrix} \right]\] As you can see...we're getting closer :)

OpenStudy (anonymous):

so all we need is to get the last 0?

hero (hero):

now simply take row 2 and subtract it from row 1 to get: \[\left[ \begin{matrix} 1 & 0 & -30 \\ 0 & 1 & 32 \end{matrix} \right]\]

OpenStudy (anonymous):

alright! now we have x=-30 y=32

hero (hero):

Yes, correct

OpenStudy (anonymous):

so how do we get to 6 and 8? the MC answers are (-6,8) (6,8) (8,-6) (6,-8)

hero (hero):

yeah, that's what I got the first time I solved this

hero (hero):

Hang on...

OpenStudy (phi):

the bottom row simplifies to 2 -3 -36

OpenStudy (phi):

start with 6x-9y=-108 6 - 9 -108 divide by 3 2 -3 -36 so your matrix is 1 1 2 2 -3 -36 now try again

hero (hero):

That's where I made the mistake...

hero (hero):

My original matrix is wrong I thought it was + 9

OpenStudy (anonymous):

okay, so we start over from there?

hero (hero):

Not to worry @HelennnnnnC I won't leave you hanging

OpenStudy (anonymous):

Thank you <3 :)

hero (hero):

I hope that link helps you out. I made sign errors while attempting to solve this time. But the correct solution is in that link above.

hero (hero):

\[\left[ \begin{matrix} -8 & -8 & -16 \\ 6 & -9 & -108 \end{matrix} \right]\] Divide the first row by 8, the second row by 3: \[\left[ \begin{matrix} 1 & 1 & 2 \\ 2 & -3 & -36 \end{matrix} \right]\] Then you could subtract R3 -2R1: \[\left[ \begin{matrix} 1 & 1 & 2 \\ 0 & -5 & -40 \end{matrix} \right]\] Then divide row 2 by -5 to get: \[\left[ \begin{matrix} 1 & 1 & 2 \\ 0 & 1 & 8 \end{matrix} \right]\] Afterwards you could subtract R1 - R2: \[\left[ \begin{matrix} 1 & 0 & -6 \\ 0 & 1 & 8 \end{matrix} \right]\]

hero (hero):

Okay, sorry...unfortunately this site doesn't make it any easier for you to correct a mistake when you type using \(LaTeX\) so it took me that long to fix it.

OpenStudy (anonymous):

thank you sooo much! <3